### v2

# Global Energy Minimization for Simplex Mesh Optimization: A Radius Ratio Approach to Sliver Elimination

Abstract
The quality of simplex mesh is crucial for the stability and accuracy of numerical simulations in finite element analysis and computational geometry. However, the presence of sliver elements in 3D simplex mesh can severely impact the results. This paper presents a novel method based on a radius ratio energy function to optimize the quality of simplex mesh elements. This method can effectively eliminate sliver elements, thereby enhancing mesh quality. The gradient of the proposed energy function can be decomposed into a matrix-vector product. With minor processing, the matrix becomes symmetric positive definite, and this symmetric positive definite matrix can serve as a preconditioner to significantly accelerate the optimization process. Experimental results demonstrate that this method has significant advantages in eliminating sliver elements and improving mesh quality.
Keywords Sliver elements  $\cdot$ Radius ratio  $\cdot$ Energy function  $\cdot$ Mesh optimization  $\cdot$ Preconditioner
1 Introduction
In numerical computation, the mesh plays an important role. Mainstream simulation methods including the Finite Element Method (FEM) and Finite Volume Method (FVM) rely fundamentally on mesh generation [12, 27]. As fundamental geometric entities, simplex find wide applications across disciplines, with simplex mesh generation techniques being studied for decades [33]. Current methods primarily fall into three categories: the Advancing-Front technique [29], spatial decomposition algorithms [37, 41], and Delaunay-based methods [1, 38].
The Delaunay algorithm has become a preferred choice for simplex mesh generation due to its algorithmic simplicity and broad applicability. However, in three-dimensional cases, it tends to produce ill-shaped tetrahedron that require optimization. Common improvement strategies include Laplacian smoothing, local transformations, and hybrid approaches [28, 20], yet these fail to eliminate persistent pathological elements known as slivers [2]. A sliver features well-shaped triangular faces but possesses near-zero volume with four vertices nearly lie in the same plane, paradoxically maintaining its Delaunay validity. Such elements severely degrade computational stability and solution accuracy.
Figure 1: Sliver element
To address the issue of sliver elements, several methods have been proposed. Cavendish [2] and Guan [15] employ local transformations and element decomposition to eliminate sliver elements. Chew [8] introduced a random point insertion optimization method to eliminate sliver elements, which was further refined by Li [23]. Engwirda [11], Cheng [7], and others have also used different point insertion techniques to eliminate sliver elements. These methods are all local optimization algorithms, which can effectively ensure the mesh quality by setting termination conditions, such as dihedral angle, radius-edge ratio, etc. [24, 34].
In the past decade, there have been a lot of research works on variational optimization methods, such as ODT [5, 3] and CVT [17, 10] algorithms. These methods define energy functions and minimize them through numerical optimization to obtain high-quality mesh. However, these algorithms still need to be processed later to eliminate sliver elements, such as perturbing slivers [39] and sliver exudation [6]. Knupp also designed a new paradigm for changing mesh node coordinates using numerical optimization, called TMOP (Target-matrix Mesh Optimization Paradigm) [21, 22]. In terms of eliminating sliver elements, Saifeng Ni [31] has also has also done related work. He used a shape matching strategy to construct an energy function, which is different from the method introduced in this paper and is applicable to tetrahedral mesh of uniform size. In this paper, we propose a new method for mesh optimization based on simplex mesh elements using radius ratio to design energy functions. We call this method is Radius-Ratio Energy function Optimization. This algorithm calculates the minimum value point of the energy function and moves the mesh nodes, which can effectively penalize sliver elements. Moreover, our algorithm inherits the ideas from ODT [4] and CVT [17], constructs a symmetric positive definite (SPD) and diagonally dominant M-matrix as a preconditioner, making it an efficient global optimization algorithm. Our algorithm can be used alone or as a post-processing of existing optimization algorithms.
The rest of this paper is organized as follows: In Section 2, we will introduce the energy function based on the radius ratio of tetrahedral elements and provide the derivation of its gradient matrix. Section 3 presents the workflow of our optimization algorithm. In Section 4, we provide several numerical examples to demonstrate the effectiveness of the algorithm. Finally, in the last section, we summarize our findings and discuss future work.
2 Radius Ratio Energy Function
The convex hull of a set of points $\{\bm{x}_{i}\}_{i=0}^{d}\subset\mathbb{R}^{d}$ that do not lie in one hyperplane
$\tau=\{\bm{x}=\sum_{i=0}^{d}\lambda_{i}\bm{x}_{i}|0\leq\lambda_{i}\leq 1,\sum_{i=0}^{d}\lambda_{i}=1\}$
is called a geometric d-simplex generated by $\{\bm{x}_{i}\}_{i=0}^{d}$. For example, an interval is a 1-simplex, a triangle is a 2-simplex, and a tetrahedron is a 3-simplex.
For a positively-oriented simplex $\tau$ in $\mathbb{R}^{d}$, $d=2,3$, the radius ratio metric is defined as follows [2]:
$\mu=\frac{R}{dr}$
where $R$ is the circumradius and $r$ is the inradius of $\tau$. $\mu\in[1,+\infty]$ With $\mu=1$ if and only if $\tau$ is equilateral simplex.
Radius ratio metric $\mu$ is a function of the coordinates of the simplex vertices. Given a triangulation $\mathcal{T}$ with $N_{v}$ vertices $\{\bm{x}_{i}\}_{i=0}^{N_{v}-1}$ and $N_{c}$ simplex $\{\tau_{n}\}_{n=0}^{N_{c}-1}$. We can define an energy function of radius ratio metric on the $\mathcal{T}$
$F=\frac{1}{N_{c}}\sum_{n=0}^{N_{c}-1}\mu_{n}$
Here we call $F$ is global radius ratio energy function. By minimizing $F$, we can develop efficient global mesh smoothing algorithm by moving the points in $\mathcal{T}$. The key problem is how to compute the gradient of $F$ about mesh point coordinates.
In this work, we present the gradient formula of $\mu_{n}$ of d-simplex $\tau_{n}$ about its every vertex. Then we design the local or global mesh smoothing algorithm to improve the mesh quality. Finally, we give some numerical tests to show the efficiency of our algorithm.
2.1 2D Case
Given a 2D triangulation $\mathcal{T}$ with $N_{v}$ vertices $\{\bm{x}_{i}\}_{i=0}^{N_{v}-1}\subset\mathbb{R}^{2}$ and $N_{c}$ triangles $\{\tau_{n}\}_{n=0}^{N_{c}-1}$. For $\forall\tau_{n}=(\bm{x}_{0},\bm{x}_{1},\bm{x}_{2})$, let $e_{0}:=(\bm{x}_{1},\bm{x}_{2}),e_{1}:=(\bm{x}_{2},\bm{x}_{0})$ and $e_{2}:=(\bm{x}_{0},\bm{x}_{1})$ be the three edges of $\tau_{n}$ with length $l_{0},l_{1}$ and $l_{2}$ respectively; $|\tau_{n}|$ be the area of $\tau_{n}$.
The circumradius of $\tau_{n}$ is:
$R_{n}=\frac{q}{4|\tau_{n}|}$
with $q=l_{0}l_{1}l_{2}$. The inradius of $\tau_{n}$ is:
$r_{n}=\frac{2|\tau_{n}|}{p}$
with $p=l_{0}+l_{1}+l_{2}$. The radius ratio metric of $\tau_{n}$ is:
$\mu_{n}=\frac{R_{n}}{2r_{n}}=\frac{pq}{16|\tau_{n}|^{2}}$
Next we compute the gradient $\mu_{n}$ about $\bm{x}_{0}$.
$\nabla_{\bm{x}_{0}}\mu_{n}~=\frac{pq}{16|\tau_{n}|^{2}}(\frac{\nabla_{\bm{x}_{0}}p}{p}+\frac{\nabla_{\bm{x}_{0}}q}{q}-\frac{2}{|\tau_{n}|}\nabla_{\bm{x}_{0}}|\tau_{n}|)$
where
$\nabla_{\bm{x}_{0}}p$ $=\frac{1}{l_{1}}(\bm{x}_{0}-\bm{x}_{2})+\frac{1}{l_{2}}(\bm{x}_{0}-\bm{x}_{1})$
$\nabla_{\bm{x}_{0}}q$ $=\frac{q}{l_{1}^{2}}(\bm{x}_{0}-\bm{x}_{2})+\frac{q}{l_{2}^{2}}(\bm{x}_{0}-\bm{x}_{1})$
$\nabla_{\bm{x}_{0}}|\tau_{n}|$ $=\frac{1}{2}\bm{W}(\bm{x}_{2}-\bm{x}_{1})$
with
$\bm{W}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$
We have
$\nabla_{\bm{x}_{0}}\mu_{n}=\mu_{n}\left\{\left[\frac{1}{pl_{1}}+\frac{1}{l_{1}^{2}}\right](\bm{x}_{0}-\bm{x}_{2})+\left[\frac{1}{pl_{2}}+\frac{1}{l_{2}^{2}}\right](\bm{x}_{0}-\bm{x}_{1})+\frac{1}{|\tau_{n}|}\bm{W}(\bm{x}_{1}-\bm{x}_{2})\right\}$
For simplicity, we denote
$c=\frac{1}{|\tau_{n}|},c_{0}=\frac{1}{pl_{0}}+\frac{1}{l_{0}^{2}},c_{1}=~\frac{1}{pl_{1}}+\frac{1}{l_{1}^{2}},c_{2}=\frac{1}{pl_{2}}+\frac{1}{l_{2}^{2}},$
$c_{01}=c_{0}+c_{1},c_{02}=c_{0}+c_{2},c_{12}=c_{1}+c_{2}$
So
$\nabla_{\bm{x}_{0}}\mu_{n}=\mu_{n}\left[c_{12}\bm{x}_{0}-(c_{2}\bm{I}-c\bm{W})\bm{x}_{1}-(c_{1}\bm{I}+c\bm{W})\bm{x}_{2}\right]$
Simiarly
$\nabla_{\bm{x}_{1}}\mu_{n}$ $=\mu_{n}\left[c_{20}\bm{x}_{1}-(c_{0}\bm{I}-c\bm{W})\bm{x}_{2}-(c_{2}\bm{I}+c\bm{W})\bm{x}_{0}\right]$
$\nabla_{\bm{x}_{2}}\mu_{n}$ $=\mu_{n}\left[c_{01}\bm{x}_{2}-(c_{1}\bm{I}-c\bm{W})\bm{x}_{0}-(c_{0}\bm{I}+c\bm{W})\bm{x}_{1}\right]$
We denote
$\bm{V}_{n}=[x_{0},x_{1},x_{2},y_{0},y_{1},y_{2}]^{T}$
Then we can write it in matrix form
$\nabla_{\bm{V}_{n}}\mu_{n}~=~\begin{bmatrix}\bm{A}_{n}~&~\bm{B}_{n}\\ -\bm{B}_{n}~&~\bm{A}_{n}\end{bmatrix}~\bm{V}_{n}$
with
$\bm{A}_{n}=\mu_{n}~\begin{pmatrix}c_{12}~&~-c_{2}~&~-c_{1}~\\ -c_{2}~&~c_{20}~&~-c_{0}~\\ -c_{1}~&~-c_{0}~&~c_{01}\end{pmatrix},\bm{B}_{n}~=~\mu_{n}\begin{pmatrix}0~&~-c~&~c~\\ c~&~0~&~-c~\\ -c~&~c~&~0\end{pmatrix}$
Matrix $\bm{A}_{n}$ is Laplacian and $\bm{B}_{n}$ is antisymmetric.
Then, we can assemble the gradient matrix of each element into a global gradient matrix. Therefore, we get the following theorem
Theorem 2.1.
For the energy function
$F=\frac{1}{N_{c}}\sum_{n=0}^{N_{c}-1}\mu_{n}$
of the triangular mesh, we have
$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}~&~\bm{B}~\\ -\bm{B}~&~\bm{A}~\\ \end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\end{bmatrix}=\bm{G}_{F}\bm{V}$
where $\bm{A}$ is $N_{v}\times N_{v}$ Laplacian matrix, and $\bm{B}$ is $N_{v}\times N_{v}$ antisymmetric matrix,
$\bm{V}=[\bm{X}^{T},\bm{Y}^{T}]^{T},\bm{X}=[x_{0},x_{1},\cdots,x_{N_{v}-1}]^{T},\bm{Y}=[y_{0},y_{1},\cdots,y_{N_{v}-1}]^{T}$
2.2 3D Case
Given a 3D Tetrahedralization $\mathcal{T}$, consisting of $N_{v}$ vertices $\{\bm{x}_{i}\}_{i=0}^{N_{v}-1}\subset\mathbb{R}^{3}$ and $N_{c}$ tetrahedrons $\{\tau_{n}\}_{n=0}^{N_{c}-1}$. For $\forall\tau_{n}=(\bm{x}_{0},\bm{x}_{1},\bm{x}_{2},\bm{x}_{3})$, let $s_{0},s_{1},s_{2},s_{3}$ be the area of triangle opposite to vertices $\bm{x}_{0},\bm{x}_{1},\bm{x}_{2},\bm{x}_{3}$; $|\tau_{n}|$ be the volume of $\tau_{n}$.
The radius of circumsphere of $\tau_{n}$ is:
$R_{n}=\frac{|\bm{d}_{0}|}{12|\tau_{n}|}$
where
$\bm{d}_{0}=\bm{v}_{30}^{2}\bm{v}_{10}\times\bm{v}_{20}+\bm{v}_{10}^{2}\bm{v}_{20}\times\bm{v}_{30}+\bm{v}_{20}^{2}\bm{v}_{30}\times\bm{v}_{10}$
with
$\bm{v}_{10}=\bm{x}_{0}-\bm{x}_{1},\bm{v}_{20}=\bm{x}_{0}-\bm{x}_{2},\bm{v}_{30}=\bm{x}_{0}-\bm{x}_{3},\bm{v}^{2}=\bm{v}\cdot\bm{v}$
The radius of inside sphere of $\tau_{n}$ is
$r_{n}~=~\frac{3|\tau_{n}|}{s}$
with $s=s_{0}+s_{1}+s_{2}+s_{3}$. The radius ratio metric of $\tau_{n}$ is
$\mu_{n}=\frac{R_{n}}{3r_{n}}=\frac{s|\bm{d}_{0}|}{108|\tau_{n}|^{2}}$
The gradient of $\mu_{n}$ at vertex $\bm{x}_{0}$ is
$\nabla_{\bm{x}_{0}}\mu_{n}=\mu_{n}\left\{\frac{1}{|\bm{d}_{0}|}\nabla_{\bm{x}_{0}}|\bm{d}_{0}|+\frac{1}{s}\nabla_{\bm{x}_{0}}s-\frac{2}{|\tau_{n}|}\nabla_{\bm{x}_{0}}|\tau_{n}|\right\}$
where
$\nabla_{\bm{x}_{0}}|\bm{d}_{0}|=$ $\frac{1}{|\bm{d}_{0}|}\{2[\bm{d}_{0}\cdot(\bm{v}_{20}\times\bm{v}_{30})\bm{v}_{10}+\bm{d}_{0}\cdot(\bm{v}_{30}\times\bm{v}_{10})\bm{v}_{20}+\bm{d}_{0}\cdot(\bm{v}_{10}\times\bm{v}_{20})\bm{v}_{30}]~$
$+\bm{d}_{0}\times[(\bm{v}_{30}^{2}-\bm{v}_{20}^{2})\bm{v}_{10}+(\bm{v}_{10}^{2}-\bm{v}_{30}^{2})\bm{v}_{20}+(\bm{v}_{20}^{2}-\bm{v}_{10}^{2})\bm{v}_{30}]\}$
$\nabla_{\bm{x}_{0}}s=(\frac{\bm{v}_{31}\cdot\bm{v}_{30}}{4s_{2}}+\frac{\bm{v}_{21}\cdot\bm{v}_{20}}{4s_{3}})\bm{v}_{10}+(\frac{\bm{v}_{32}\cdot\bm{v}_{30}}{4s_{1}}+\frac{\bm{v}_{12}\cdot\bm{v}_{10}}{4s_{3}})\bm{v}_{20}+(\frac{\bm{v}_{23}\cdot\bm{v}_{20}}{4s_{1}}+\frac{\bm{v}_{13}\cdot\bm{v}_{10}}{4s_{2}})\bm{v}_{30}$
$\nabla_{\bm{x}_{0}}|\tau_{n}|=-\frac{1}{18}\left\{[\bm{x}_{2}-2\bm{x}_{3}+\bm{x}_{1}]_{\times}\bm{v}_{10}+[\bm{x}_{3}-2\bm{x}_{1}+\bm{x}_{2}]_{\times}\bm{v}_{20}+[\bm{x}_{1}-2\bm{x}_{2}+\bm{x}_{3}]_{\times}\bm{v}_{30}\right\}$
We denote
$\bm{V}_{n}=[x_{0},x_{1},x_{2},x_{3},y_{0},y_{1},y_{2},y_{3},z_{0},z_{1},z_{2},z_{3}]^{T}$
and calculate the matrix form of $\nabla_{\bm{x}}|\bm{d}_{0}|,\nabla_{\bm{x}}s,\nabla_{\bm{x}}|\tau_{n}|$.
For $\nabla_{\bm{x}}|\bm{d}_{0}|$, we denote
$c_{23}=\bm{d}_{0}\cdot(\bm{v}_{20}\times\bm{v}_{30}),c_{31}=\bm{d}_{0}\cdot(\bm{v}_{30}\times\bm{v}_{10}),c_{12}=\bm{d}_{0}\cdot(\bm{v}_{10}\times\bm{v}_{20})$
$k_{23}=\bm{v}_{30}^{2}-\bm{v}_{20}^{2},k_{31}=\bm{v}_{10}^{2}-\bm{v}_{30}^{2},k_{12}=\bm{v}_{20}^{2}-\bm{v}_{10}^{2}$
$[\bm{d}_{0}]_{\times}=\bm{D}_{0}=\begin{bmatrix}0~&~-d_{2}~&~d_{1}~\\ d_{2}~&~0~&~-d_{0}~\\ -d_{1}~&~d_{0}~&~0\end{bmatrix}$
so we have
$\nabla_{\bm{x}_{0}}|\bm{d}_{0}|=\frac{1}{12|\tau_{n}|R_{n}}\{[2(c_{23}+c_{31}+c_{12})+(k_{23}+k_{31}+k_{12})\bm{D}_{0}]\bm{x}_{0}-(2c_{23}+k_{23}\bm{D}_{0})\bm{x}_{1}-(2c_{31}+k_{31}\bm{D}_{0})\bm{x}_{2}-(2c_{12}+k_{12}\bm{D}_{0})\bm{x}_{3}\}$
$\nabla_{\bm{x}_{1}}|\bm{d}_{0}|$ $=\frac{1}{12|\tau_{n}|R_{n}}\{(-2c_{23}+k_{23}\bm{D}_{0})\bm{x}_{0}+2c_{23}\bm{x}_{1}-\bm{v}_{30}^{2}\bm{D}_{0}\bm{x}_{2}+\bm{v}_{20}^{2}\bm{D}_{0}\bm{x}_{3}\}$
$\nabla_{\bm{x}_{2}}|\bm{d}_{0}|$ $=\frac{1}{12|\tau_{n}|R_{n}}\{(-2c_{31}+k_{31}\bm{D}_{0})\bm{x}_{0}+\bm{v}_{30}^{2}\bm{D}_{0}\bm{x}_{1}+2c_{31}\bm{x}_{2}-\bm{v}_{10}^{2}\bm{D}_{0}\bm{x}_{3}\}$
$\nabla_{\bm{x}_{3}}|\bm{d}_{0}|$ $=\frac{1}{12|\tau_{n}|R_{n}}\{(-2c_{12}+k_{12}\bm{D}_{0})\bm{x}_{0}-\bm{v}_{20}^{2}\bm{D}_{0}\bm{x}_{1}+\bm{v}_{10}^{2}\bm{D}_{0}\bm{x}_{2}+2c_{12}\bm{x}_{3}\}$
Then we can write $\nabla_{\bm{V}_{n}}|\bm{d}_{0}|$ in matrix form
$\nabla_{\bm{V}_{n}}|\bm{d}_{0}|~=~\frac{1}{12|\tau_{n}|R_{n}}\begin{bmatrix}\bm{M}~&~-d_{2}\bm{K}~&~d_{1}\bm{K}~\\ d_{2}\bm{K}~&~\bm{M}~&~-d_{0}\bm{K}~\\ -d_{1}\bm{K}~&~d_{0}\bm{K}~&~\bm{M}\end{bmatrix}~\bm{V}_{n}$
with
$\bm{M}=\begin{bmatrix}2c~&~-2c_{23}~&~-2c_{31}~&~-2c_{12}~\\ -2c_{23}~&~2c_{23}~&~0~&~0~\\ -2c_{31}~&~0~&~2c_{31}~&~0~\\ -2c_{12}~&~0~&~0~&~2c_{12}\end{bmatrix},\bm{K}=\begin{bmatrix}0~&~-k_{23}~&~-k_{31}~&~-k_{12}~\\ k_{23}~&~0~&~-\bm{v}_{30}^{2}~&~\bm{v}_{20}^{2}~\\ k_{31}~&~\bm{v}_{30}^{2}~&~0~&~-\bm{v}_{10}^{2}~\\ k_{12}~&~-\bm{v}_{20}^{2}~&~\bm{v}_{10}^{2}~&~0\end{bmatrix}$
$c=c_{23}+c_{31}+c_{12}$, matrix $\bm{M}$ is symmetric, $\bm{K}$ is antisymmetric.
For $\nabla_{\bm{x}}s$, we denote
$p_{0}$ $=~\frac{\bm{v}_{31}^{2}}{4s_{2}}+\frac{\bm{v}_{21}^{2}}{4s_{3}}+\frac{\bm{v}_{32}^{2}}{4s_{1}},\quad~p_{1}~=~\frac{\bm{v}_{32}^{2}}{4s_{0}}+\frac{\bm{v}_{02}^{2}}{4s_{3}}+\frac{\bm{v}_{30}^{2}}{4s_{2}},\quad$
$~p_{2}$ $=~\frac{\bm{v}_{30}^{2}}{4s_{1}}+\frac{\bm{v}_{10}^{2}}{4s_{3}}+\frac{\bm{v}_{31}^{2}}{4s_{0}},\quad p_{3}~=~\frac{\bm{v}_{10}^{2}}{4s_{2}}+\frac{\bm{v}_{20}^{2}}{4s_{1}}+\frac{\bm{v}_{21}^{2}}{4s_{0}},\quad$
$q_{01}$ $=-(\frac{\bm{v}_{31}\cdot\bm{v}_{30}}{4s_{2}}+\frac{\bm{v}_{21}\cdot\bm{v}_{20}}{4s_{3}}),q_{02}=-(\frac{\bm{v}_{32}\cdot\bm{v}_{30}}{4s_{1}}+\frac{\bm{v}_{12}\cdot\bm{v}_{10}}{4s_{3}}),q_{03}=-(\frac{\bm{v}_{23}\cdot\bm{v}_{20}}{4s_{1}}+\frac{\bm{v}_{13}\cdot\bm{v}_{10}}{4s_{2}}),$
$q_{12}$ $=-(\frac{\bm{v}_{32}\cdot\bm{v}_{31}}{4s_{0}}+\frac{\bm{v}_{02}\cdot\bm{v}_{01}}{4s_{3}}),q_{13}=-(\frac{\bm{v}_{03}\cdot\bm{v}_{01}}{4s_{2}}+\frac{\bm{v}_{23}\cdot\bm{v}_{21}}{4s_{0}}),q_{23}=-(\frac{\bm{v}_{13}\cdot\bm{v}_{12}}{4s_{0}}+\frac{\bm{v}_{03}\cdot\bm{v}_{02}}{4s_{1}})$
We have
$\nabla_{\bm{x_{0}}}s$ $=p_{0}\bm{x}_{0}+q_{01}\bm{x}_{1}+q_{02}\bm{x}_{2}+q_{03}\bm{x}_{03},\nabla_{\bm{x}_{1}}s=q_{01}\bm{x}_{0}+p_{1}\bm{x}_{1}+q_{12}\bm{x}_{2}+q_{23}\bm{x}_{3}$
$\nabla_{\bm{x}_{2}}s$ $=q_{02}\bm{x}_{0}+q_{12}\bm{x}_{1}+p_{2}\bm{x}_{2}+q_{23}\bm{x}_{3},\nabla_{\bm{x}_{3}}s=q_{03}\bm{x}_{0}+q_{13}\bm{x}_{1}+q_{23}\bm{x}_{2}+p_{3}\bm{x}_{3}$
Then we can write $\nabla_{\bm{V}_{n}}s$ in matrix form
$\nabla_{\bm{V}_{n}}\mu_{n}~=~\begin{bmatrix}\bm{S}~&~~&~~\\ ~&~\bm{S}~&~~\\ ~&~~&~\bm{S}\end{bmatrix}~\bm{V}_{n}$
with
$\bm{S}~=~\begin{bmatrix}p_{0}&q_{01}&q_{02}&q_{03}\\ q_{01}&p_{1}&~q_{12}&~q_{13}\\ q_{02}&~q_{12}&p_{2}&q_{23}\\ q_{03}&q_{13}&q_{23}&p_{3}\end{bmatrix}$
Matrix $\bm{S}$ is symmetric.
For $\nabla_{\bm{x}}|\tau_{n}|$, we have
$\nabla_{\bm{x}_{0}}|\tau_{n}|$ $=\frac{1}{6}(\bm{x}_{2}\times\bm{x}_{1}+\bm{x}_{3}\times\bm{x}_{2}+\bm{x}_{1}\times\bm{x}_{3}),\nabla_{\bm{x}_{1}}|\tau_{n}|=\frac{1}{6}(\bm{x}_{3}\times\bm{x}_{0}+\bm{x}_{0}\times\bm{x}_{2}+\bm{x}_{2}\times\bm{x}_{3}),$
$\nabla_{\bm{x}_{2}}|\tau_{n}|$ $=\frac{1}{6}(\bm{x}_{1}\times\bm{x}_{0}+\bm{x}_{3}\times\bm{x}_{1}+\bm{x}_{0}\times\bm{x}_{3}),\nabla_{\bm{x}_{3}}|\tau_{n}|=\frac{1}{6}(\bm{x}_{2}\times\bm{x}_{0}+\bm{x}_{0}\times\bm{x}_{1}+\bm{x}_{1}\times\bm{x}_{2})$
We denote
$\bm{C}_{0}=\begin{bmatrix}0~&~x_{2}~&~x_{3}~&~x_{1}~\\ x_{3}~&~0~&~x_{0}~&~x_{2}~\\ x_{1}~&~x_{3}~&~0~&~x_{0}~\\ x_{2}~&~x_{0}~&~x_{1}~&0\end{bmatrix},\bm{C}_{1}=\begin{bmatrix}0~&~y_{2}~&~y_{3}~&~y_{1}~\\ y_{3}~&~0~&~y_{0}~&~y_{2}~\\ y_{1}~&~y_{3}~&~0~&~y_{0}~\\ y_{2}~&~y_{0}~&~y_{1}~&0\end{bmatrix},\bm{C}_{2}=\begin{bmatrix}0~&~z_{2}~&~z_{3}~&~z_{1}~\\ z_{3}~&~0~&~z_{0}~&~z_{2}~\\ z_{1}~&~z_{3}~&~0~&~z_{i}~\\ z_{2}~&~z_{0}~&~z_{1}~&0\end{bmatrix}$
Then we have
$\nabla_{\bm{V}_{n}}|\tau_{n}|=\frac{1}{6}\begin{bmatrix}0~&~-\bm{C}_{2}~&~\bm{C}_{1}~\\ \bm{C}_{2}~&~0~&~-\bm{C}_{0}~\\ -\bm{C}_{1}~&~\bm{C}_{0}~&~0\end{bmatrix}\bm{V}_{n}=\bm{C}\bm{V}_{n}$
Where $|\tau_{n}|$ is a cubic form with respect to $\bm{V}_{n}$, we have the following lemma for cubic forms.
Lemma 2.2.
If the function $f$ is a cubic form with respect to $\bm{X}=(x_{0},x_{1},\cdots,x_{n})$, then there exist matrix $\bm{C}$ and $\bm{E}=\frac{1}{2}(\bm{C}+\bm{C}^{T})$ such that
$\nabla f=\bm{C}\bm{X}=\bm{E}\bm{X}$
Proof.
If the function $f$ is a cubic form with respect to $\bm{X}=(x_{0},x_{1},\cdots,x_{n})$, then there exists a third-order matrix $\bm{U}$ such that
$f~=~U_{ijk}x_{i}x_{j}x_{k}$
Let $W_{ik}^{j}=\frac{1}{2}(U_{ijk}+U_{kji})$, for any fixed $j$, $\bm{W}^{j}$ is a symmetric matrix, and we have
$f~=W_{ik}^{j}~x_{i}~x_{j}x_{k}$
So we have $\nabla_{x_{k}}f=U_{ijk}x_{i}x_{j}=W_{ik}^{j}x_{i}x_{j}$, let $C_{ki}~=~U_{ijk}x_{j},~E_{ki}~=~W_{ik}^{j}x_{j}$, so $\bm{E}~=~\frac{1}{2}(\bm{C}+\bm{C}^{T})$, then we can express $\nabla f$ as
$\nabla f=\bm{C}\bm{X}=\bm{E}\bm{X}$
From the above theorem, we can get $\nabla_{\bm{V}_{n}}|\tau_{n}|=\frac{1}{2}(\bm{C}+\bm{C}^{T})\bm{V}_{n}$.
$\frac{1}{2}(\bm{C}+\bm{C}^{T})=\frac{1}{12}\begin{bmatrix}0~&~\bm{C}_{2}^{T}-\bm{C}_{2}~&~\bm{C}_{1}^{T}-\bm{C}_{1}~\\ \bm{C}_{2}-\bm{C}_{2}^{T}~&~0~&~\bm{C}_{0}^{T}-\bm{C}_{0}~\\ \bm{C}_{1}^{T}-\bm{C}_{1}~&~\bm{C}_{0}^{T}-\bm{C}_{0}~&~0\end{bmatrix}$
It is evident that each block matrix in $\frac{1}{2}(\bm{C}+\bm{C}^{T})$ is an antisymmetric matrix.
From the equation
$\nabla_{\bm{x}_{0}}\mu_{n}=\mu_{n}\left\{\frac{1}{|\bm{d}_{0}|}\nabla_{\bm{x}_{0}}|\bm{d}_{0}|+\frac{1}{s}\nabla_{\bm{x}_{0}}s-\frac{2}{|\tau_{n}|}\nabla_{\bm{x}_{0}}|\tau_{n}|\right\}$
and the gradients $\nabla_{\bm{V}_{n}}|\bm{d}_{0}|$, $\nabla_{\bm{V}_{n}}s$, and $\nabla_{\bm{V}_{n}}|\tau_{n}|$, we can obtain
$\nabla_{\bm{V}_{n}}\mu_{n}=\mu_{n}\begin{bmatrix}\bm{A}_{n}~&~\bm{B}_{0n}~&~\bm{B}_{1n}~\\ -\bm{B}_{2n}~&~\bm{A}_{n}~&~\bm{B}_{0n}~\\ -\bm{B}_{1n}~&~-\bm{B}_{0n}~&~\bm{A}_{n}\end{bmatrix}\bm{V}_{n}$
with
$\bm{A}_{n}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}+\frac{1}{s}\bm{S},~\bm{B}_{0n}=\frac{-d_{0}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{0}^{T}-\bm{C}_{0}),$
$\bm{B}_{1n}=\frac{d_{1}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{1}^{T}-\bm{C}_{1}),\bm{B}_{2n}=\frac{-d_{2}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{2}^{T}-\bm{C}_{2})$
Obviously, $\bm{A}_{n}$ is symmetric matrix, and $\bm{B}_{0n}$, $\bm{B}_{1n}$, $\bm{B}_{2n}$ are antisymmetric matrices.
Then, we can assemble the gradient matrix of each element into a global gradient matrix. Therefore, we get the following theorem
Theorem 2.3.
For the energy function
$F=\frac{1}{N_{c}}\sum_{n=0}^{N_{c}-1}\mu_{n}$
of the tetrahedral mesh,we have
$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}~&~\bm{B}_{2}~&~\bm{B}_{1}~\\ -\bm{B}_{2}~&~\bm{A}~&~\bm{B}_{0}~\\ -\bm{B}_{1}~&~-\bm{B}_{0}~&~\bm{A}\end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\\ \bm{Z}\end{bmatrix}=\bm{G}_{F}\bm{V}~$
where $\bm{A}$ is $N_{v}\times N_{v}$ symmetric matrix, and $\bm{B}_{0}$, $\bm{B}_{1}$, $\bm{B}_{2}$ are $N_{v}\times N_{v}$ antisymmetric matrices.
$\bm{V}=[\bm{X}^{T},\bm{Y}^{T},\bm{Z}^{T}]^{T},\bm{X}=[x_{0},x_{1},\cdots,x_{N_{v}-1}]^{T},\bm{Y}=[y_{0},y_{1},\cdots,y_{N_{v}-1}]^{T},\bm{Z}=[z_{0},z_{1},\cdots,z_{N_{v}-1}]^{T}$
3 Energy Optimization
In Chapter 2, we formulate a radius ratio energy function $F(\bm{x})$, where $\bm{x}$ represents the coordinate vector of mesh nodes. To obtain the optimal mesh configuration under the current connectivity, we need to find the minimizer of $F(\bm{x})$. We derive the analytical gradient $\nabla F(\bm{x})$, which enables the implementation of various optimization algorithms such as quasi-Newton methods and nonlinear conjugate gradient (NLCG) algorithms. Notably, the gradient can be expressed in matrix-vector product form:
$\nabla F(\bm{x})=\bm{G}_{F}\bm{x}$
By constructing a preconditioner derived from $\bm{G}_{F}$, we achieve significant improvements in the computational efficiency of the optimization algorithm.
$\bm{G}_{F}$ is a block matrix, with its diagonal block $\bm{A}$ formed through finite element assembly of per-cell gradient diagonal matrix $\bm{A}_{n}$
$\bm{A}_{n}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}+\frac{1}{s}\bm{S}$
$\bm{S}$ is laplacian,
$\bm{M}=\begin{bmatrix}2c~&~-2c_{23}~&~-2c_{31}~&~-2c_{12}~\\ -2c_{23}~&~2c_{23}~&~0~&~0~\\ -2c_{31}~&~0~&~2c_{31}~&~0~\\ -2c_{12}~&~0~&~0~&~2c_{12}\end{bmatrix}$
$c=c_{12}+c_{23}+c_{31}$. However, since the terms $c_{12},c_{23},c_{31}$ may take negative values, we define
$\bm{M}^{abs}=\begin{bmatrix}2|c_{12}|+|c_{23}|+|c_{31}|~&~-2|c_{23}|~&~-2|c_{31}|~&~-2|c_{12}|~\\ -2|c_{23}|~&~2|c_{23}|~&~0~&~0~\\ -2|c_{31}|~&~0~&~2|c_{31}|~&~0~\\ -2|c_{12}|~&~0~&~0~&~2|c_{12}|\end{bmatrix}$
$\bm{A}_{n}^{abs}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}^{abs}+\frac{1}{s}\bm{S}$
By this way, the matrix $\bm{A}^{abs}_{n}$ is symmetric positive semi-definite, and consequently, the assembled global matrix $\bm{A}_{abs}$ inherits this symmetry and semi-definiteness. During optimization, by fixing a subset or all boundary nodes and removing the corresponding rows and columns from $\bm{A}_{abs}$, we obtain the reduced matrix $\bm{P}$. For the matrix $\bm{P}$, we have the following theorem:
Theorem 3.1.
If the region where the mesh is generated is a connected region, then the matrix $\bm{P}$ is a symmetric positive definite matrix.
Proof.
According to the known conditions, we have the following inferences:
  1. 1.
    Since $\bm{P}$ is obtained by deleting some rows and columns corresponding to boundary points from $\bm{P}$, $\bm{P}$ has at least one row that satisfies strict diagonal dominance. Therefore, it is a weakly diagonally dominant matrix with an eigenvalue greater than or equal to 0.
  2. 2.
    The directed graph corresponding to $\bm{P}$ represents the node connectivity topology of the grid. If the region where the mesh is generated is a connected region, then $\bm{P}$ is irreducible.
  3. 3.
    $\bm{P}$ is an irreducible weakly diagonally dominant matrix, so $\bm{P}$ is non-singular [19] and has non-zero eigenvalues.
Combining these results,the eigenvalue of $\bm{P}$ is greater than 0, and $\bm{P}$ is obviously a symmetric matrix, so $\bm{P}$ is a symmetric positive definite matrix. Furthermore, fixing at least one boundary node ensures $\bm{P}$ remains symmetric positive definite. ∎
Due to the favorable properties of the matrix $\bm{P}$, we employ it as the preconditioner.
3.1 Fixed-point iteration algorithm
Given the nodal coordinate vector $\bm{V}$ and the gradient matrix
$\bm{G}_{F}=\begin{bmatrix}\bm{A}~&~\bm{B}_{2}~&~\bm{B}_{1}~\\ -\bm{B}_{2}~&~\bm{A}~&~\bm{B}_{0}~\\ -\bm{B}_{1}~&~-\bm{B}_{0}~&~\bm{A}\end{bmatrix}$
the optimization can be solved using the fixed-point iteration method.
The gradient condition $\bm{G}_{F}\bm{V}=0$ can be formulated as a system of equations
$\begin{cases}\bm{A}\bm{X}+\bm{B}_{2}\bm{Y}+\bm{B}_{1}\bm{Z}=\bm{0}~\\ -\bm{B}_{2}\bm{X}+\bm{A}\bm{Y}+\bm{B}_{0}\bm{Z}=\bm{0}~\\ -\bm{B}_{1}\bm{X}-\bm{B}_{0}\bm{Y}+\bm{A}\bm{Z}=\bm{0}\end{cases}$
By employing the fixed-point iteration method for optimization with fixed boundaries and decomposing $\bm{V}=\bm{V}_{free}+\bm{V}_{fix}$, the system of equations can be expressed as
$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}_{2}\bm{Y}^{old}-\bm{B}_{1}\bm{Z}^{old}-\bm{A}\bm{X}^{old}_{fix}~\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}_{2}\bm{X}^{old}-\bm{B}_{0}\bm{Z}^{old}-\bm{A}\bm{Y}^{old}_{fix}~\\ \bm{A}\bm{Z}_{free}^{new}=\bm{B}_{1}\bm{X}^{old}+\bm{B}_{0}\bm{Y}^{old}-\bm{A}\bm{Z}^{old}_{fix}\end{cases}$
In practical implementations, substituting $\bm{A}$ with $\bm{A}_{abs}$ achieves enhanced computational performance.
For the two-dimensional case, the system of equations is formulated as
$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}\bm{Y}^{old}-\bm{A}\bm{X}^{old}_{fix}~\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}\bm{X}^{old}-\bm{A}\bm{Y}^{old}_{fix}\end{cases}$
Let $d=\bm{V}^{new}-\bm{V}^{old}$ denote the search direction, and update the mesh nodal coordinate vector as
$\bm{V}^{new}=\bm{V}+\lambda d$
In 2D cases, set $\lambda=1$ with backtracking for step size control; for 3D cases, set $\lambda=1$ with the strong Wolfe conditions for step size regulation.
3.2 PLBFGS Algorithm
Quasi-Newton methods approximate the Hessian matrix using gradient information of the objective function, and one of the most popular quasi-Newton algorithm is the BFGS method. We describe one iteration of the preconditioned limited memory version of the BFGS (PLBFGS), for more details, please refer to the  [32, 25].
First, we initialize
$\delta=\begin{cases}0,\quad\text{if}~{}k\leq~m~\\ k-m,\quad\text{if}~{}k>m\end{cases};L=\begin{cases}k,\quad\text{if}~{}k\leq~m~\\ m,\quad\text{if}~{}k>m\end{cases};\bm{q}_{L}=\nabla F(\bm{x}_{k})$
Each iteration of the PLBFGS algorithm is as follows:
Algorithm 1 PLBFGS Algorithm
1:  First Loop:
2:  for $i=L-1,L-2,\cdots,1,0$do
3:     $j=i+\delta$
4:     $\alpha_{i}=\rho_{j}\bm{s}_{j}^{T}\bm{q}_{i+1}$
5:     $\bm{q}_{i}=\bm{q}_{i+1}-\alpha_{i}\bm{y}_{j}$
6:  end for
7:  Set search direction: $\bm{P}_{k}\bm{r}_{0}=\bm{q}_{0}$(if no preconditioner,it is $\bm{r}_{0}=\bm{q}_{0}$)
8:  Second Loop:
9:  for $i=0,1,2,\cdots,L-2,L-1$do
10:     $j=i+\delta$
11:     $\beta_{j}=\rho_{j}\bm{y}_{j}^{T}\bm{r}_{i}$
12:     $\bm{r}_{i+1}=\bm{r}_{i}+(\alpha_{i}-\beta_{i})\bm{s}_{j}$
13:  end for
14:  $d_{k}=\bm{r}_{L}$
15:  $\bm{x}_{k+1}=\bm{x}_{k}+\lambda_{k}d_{k}$
The matrix $\bm{P}_{k}$ is a chosen preconditioner, $\lambda_{k}$ is given by the line search satisfying the strong Wolfe conditions [32] with an initial guess $\lambda_{k}=1$, and
$\bm{s}_{k}=\bm{s}_{k+1}-\bm{s}_{k},\bm{y}_{k}=\nabla F(\bm{x}_{k+1})-\nabla F(\bm{x}_{k}),\rho_{k}=\frac{1}{\bm{y}_{k}^{T}\bm{s}_{k}}$
Since our preconditioner is symmetric positive definite, we employ the conjugate gradient (CG) algorithm [30] to compute $\bm{P}_{k}\bm{r}_{0}=\bm{q}_{0}$. For large-scale problems, the algebraic multigrid (AMG) method [26, 36] can be substituted to enhance computational efficiency.
3.3 PNLCG Algorithm
The Nonlinear Conjugate Gradient method(NLCG) [32, 16] is an iterative algorithm for solving nonlinear optimization problems. It extends the Conjugate Gradient (CG) method.
Given an initial value $\bm{x}_{0}$ and initial gradient $\bm{g}_{0}=\nabla F(\bm{x}_{0})$, set the search direction $\bm{p}_{0}=-\bm{g}_{0}$. Then, one iteration step of the NLCG method proceeds as follows
Algorithm 2 NLCG Algorithm
1:  Update $\bm{x}_{k+1}=\bm{x}_{k}+\lambda_{k}\bm{p}_{k}$
2:  Set $\bm{g}_{k+1}=\nabla F(\bm{x}_{k+1})$
3:  Calculate Scalar coefficient $\beta_{k+1}$
4:  Update search direction $\bm{p}_{k+1}=-\bm{g}_{k+1}+\beta_{k+1}\bm{p}_{k}$
Given a symmetric positive definite matrix $\bm{P}_{0}$ as a preconditioner, solve $\bm{P}_{0}\hat{\bm{g}}_{0}=\bm{g}_{0}$ for the preconditioned gradient $\hat{\bm{g}}_{0}$, and $\bm{P}_{0}\bm{p}_{0}=-\bm{g}_{0}$ for the search direction $\bm{p}_{0}$. Since $\bm{P}_{0}$ is symmetric positive definite, we solve these systems using the conjugate gradient method. One iteration step of the PNLCG method proceeds as follows:
Algorithm 3 PNLCG Algorithm
1:  Update $\bm{x}_{k+1}=\bm{x}_{k}+\lambda_{k}\bm{p}_{k}$
2:  Update $\bm{P}_{k+1}$,calculate $\bm{P}_{k+1}\hat{\bm{g}}_{k+1}=-\bm{g}_{k+1}$
3:  Calculate Scalar coefficient $\beta_{k+1}$
4:  Update search direction $\bm{p}_{k+1}=-\bm{g}_{k+1}+\beta_{k+1}\bm{p}_{k}$
$\lambda_{k}$ is given by line search satisfying the strong Wolfe conditions with an initial guess $\lambda_{k}=1$. The scalar coefficient $\beta_{k}$ has multiple calculation formulas, see [16]. Notable $\beta_{k}$ formulas include: Polak-Ribière ($\beta_{k}^{PR}$[35], Hestenes–Stiefel ($\beta_{k}^{HS}$[18], Fletcher–Reeves ($\beta_{k}^{FR})$ [13], and Dai-Yuan($\beta_{k}^{DY}$[9]. We adopt the Polak-Ribière $\beta_{k}$ formula, defined as:
$\beta_{k+1}^{PR}=\frac{\bm{g}_{k+1}^{T}(\bm{g}_{k+1}-\bm{g}_{k})}{\bm{g}_{k}^{T}\bm{g}_{k}}$
4 Experiment and Comparisons
We implemented the algorithm in Python and integrated it into FEALPy  [40], an open-source numerical PDE solver package. To demonstrate its optimization capability for sliver elements, we also developed a Python implementation of the ODT local mesh optimization method for comparison. Section 4.1 presents 2D triangular mesh examples, while Section 4.2 provides 3D tetrahedral mesh examples. The radius ratio $\mu=\frac{R}{dr}$ of inscribed to circumscribed spheres is adopted as the quality metric.
We run all experiments on a laptop running Ubuntu 22.04. The CPU is an AMD Ryzen7 4800H(2.9GHz,8MB cache) and 16GB DDR4 RAM(3200MHz).
4.1 2D TriangleMesh
First, we present a specific example to demonstrate the effectiveness of our global algorithm. We perturb the positions of two vertices in a uniform mesh of equilateral triangles, refine it(Fig. 2(a)), and then apply the ODT algorithm combined with the radius ratio optimization. The latter employs a fixed-point iteration algorithm with a step size set to 1 and controlled via backtracking, similar experiment is already in Chen & Holst [4].
Figure 2: Comparison of mesh obtained by local and global methods
It can be observed that the local ODT algorithms improved the mesh quality to some extent after 100 iterations. However, even with further iterations, it is difficult to achieve significant improvement in mesh quality. In contrast, the global algorithm only required 13 iterations and achieved a globally optimal mesh. In this particular case, the mesh topology is already optimal, which is why the global algorithm reached the global optimum. In general cases, achieving global optimality is not guaranteed, but we can still observe the efficiency of our global algorithm. In fact, our algorithm theoretically performs very well in terms of optimization when the topology is already optimal.
Now we present a more general example. We first utilize the open-source mesh generation software Gmsh [14] to construct the geometric domain and generate the initial mesh, then compare the performance of several optimization algorithms for mesh improvement.
Figure 3: Comparison of our mesh optimization method with ODT
As shown in the Fig. 3, in the two-dimensional case, our optimization effect is very good, and we use fewer optimization times to get better results than ODT optimization algorithms.
4.2 3D TetrahedronMesh
For tetrahedral mesh, we employ the PLBFGS and PNLCG methods for optimization. Both PLBFGS algorithm and PNLCG algorithm are implemented in FEALPy. Since our algorithm performs node relocation without altering topology, we combine it with the ODT algorithm, which effectively improves mesh topology and enhances the optimized mesh quality.
Figure 4: Sphere example
Comparative analysis reveals that the radius ratio optimization algorithm significantly improves the quality of the worst elements, while the ODT algorithm enhances overall mesh quality but offers limited gains for poor-quality elements. Combining ODT with radius ratio optimization achieves both global refinement and targeted improvement of suboptimal elements. We present the results of 15 and 75 ODT iterations. After 75 ODT iterations, the mesh quality is difficult to improve. At this time, a small number of radius ratio mesh optimization iterations can greatly improve the mesh quality. Similarly, even limited ODT iterations followed by radius ratio optimization outperform standalone use of either method.
Additionally, the examples compare scenarios with and without preconditioners. We report the number of optimization iterations and Wolfe line search steps (in parentheses). The results demonstrate that preconditioning significantly reduces both iteration counts and search steps(often by nearly half or more) and takes the less time, while achieving comparable or superior optimization outcomes to non-preconditioned cases.
Figure 5: For the L-shaped domain, the corner points and points on the domain boundary edges are fixed, while points on the boundary faces are allowed to move along the tangential directions of the faces. ODT optimization converges after 36 iterations.
Figure 6: 12 intersecting spheres example
Table 1: Comparison of different methods and different numbers of iterations
Model Method Iterations times Min radius ratio <0.3 Number of tets Time(sec.)
Sphere Init / 0.01317 391 19784 /
ODT 15 0.05079 106 19072 4.18
ODT 75 0.06376 58 18852 20.31
LBFGS 17(85) 0.10295 53 19784 7.95
PLBFGS 13(54) 0.18480 44 19784 5.84
NLCG 30(107) 0.16996 45 19784 9.89
PNLCG 16(69) 0.17120 43 19784 6.97
ODT+LBFGS 15+15(59) 0.28190 1 19072 4.18+3.64
ODT+PLBFGS 15+7(29) 0.30972 0 19072 4.18+2.16
ODT+PNLCG 15+7(38) 0.29942 1 19072 4.18+2.65
ODT+LBFGS 75+13(33) 0.31686 0 18852 20.31+2.04
ODT+PLBFGS 75+7(27) 0.32546 0 18852 20.31+2.01
ODT+PNLCG 75+8(34) 0.32572 0 18852 20.31+2.45
Lshape Init / 0.00971 610 32827 /
ODT 15 0.01938 225 31846 13.48
ODT 36 0.06550 96 31371 31.49
LBFGS 14(109) 0.00556 6 32827 13.49
PLBFGS 24(145) 0.20427 10 32827 27.34
NLCG 10(82) 0.01216 12 32827 14.06
PNLCG 16(117) 0.12828 5 32827 22.22
ODT+LBFGS 15+22(91) 0.24053 4 31846 13.48+12.07
ODT+PLBFGS 15+12(61) 0.22503 4 31846 13.48+9.26
ODT+PNLCG 15+14(76) 0.19577 4 31846 13.48+11.50
ODT+LBFGS 36+12(51) 0.29676 1 31371 22.22+6.72
ODT+PLBFGS 36+8(32) 0.34373 0 31371 22.22+5.08
ODT+PNLCG 36+9(51) 0.31098 0 31371 22.22+7.57
12 spheres
intersecting
init / 0.00408 1470 70105 /
ODT 15 0.01119 843 68962 30.35
ODT 75 0.08138 188 66581 144.77
LBFGS 48(224) 0.17128 215 70105 83.62
PLBFGS 25(139) 0.19374 219 70105 56.42
NLCG 37(188) 0.09946 235 70105 36.26
PNLCG 16(88) 0.14231 217 70105 36.26
ODT+LBFGS 15+21(107) 0.22678 78 68962 30.35+30.85
ODT+PLBFGS 15+12(54) 0.21974 68 68962 30.35+17.92
ODT+PNLCG 15+11(61) 0.23623 66 68692 30.35+19.89
ODT+LBFGS 75+13(56) 0.24514 3 66581 144.77+15.74
ODT+PLBFGS 75+6(10) 0.28008 2 66581 144.77+4.39
ODT+PNLCG 75+7(35) 0.25638 3 66581 144.77+11.7
Table  1 presents quality statistics and efficiency comparisons for all mesh models, with iteration counts in parentheses indicating Wolfe line search steps. The data demonstrate that preconditioned optimization outperforms non-preconditioned methods in both efficiency and results. Initial mesh quality significantly influences optimization outcomes: increased ODT optimization iterations improve both effectiveness and efficiency of the radius-ratio energy optimization. Nevertheless, even with limited ODT iterations, our algorithm achieves substantial improvements.
5 Conclusion and further work
In this paper, we propose a radius-ratio energy-based global mesh optimization algorithm that effectively suppresses sliver elements and improves tetrahedral mesh quality. We also design an efficient preconditioner to substantially enhance optimization efficiency. Our algorithm can operate independently or synergistically with existing mesh optimization methods. Benefiting from the preconditioner, it significantly improves mesh quality with minimal computational overhead when applied as a post-processing step to other optimization techniques.
Our algorithm is a vertex smoothing algorithm that does not include topological optimization. It performs well when mesh topology is already good. When our algorithm is combined with ODT algorithm, it can achieve good optimization results because ODT algorithm includes topology changes and effectively optimizes the mesh topology. When used alone, our algorithm can also optimize sliver elements, although the results are less effective than when combined with ODT. Furthermore, for surface boundaries, we currently use a fixed boundary point method, which also limits the optimization effect. Therefore, in future works, we plan to develop more effective boundary processing solution and add an edge/face exchange algorithm to further improve the algorithm, enhance the reliability of the algorithm, and extend this mesh optimization algorithm to anisotropic mesh. We also hope to optimize larger-scale mesh and further improve computational efficiency by using GPU parallel technology.
Acknowledgement
This work was supported in part by the National Key R&D Program of China (2024YFA1012600), the National Natural Science Foundation of China (NSFC) (Grant No. 12371410, 12261131501), and the Graduate innovation Project of Xiangtan University(Grant No.XDCX2024Y183)

### v3

@@
 Next we compute the gradient $\mu_{n}$ about $\bm{x}_{0}$.
-$\nabla_{\bm{x}_{0}}\mu_{n}~=\frac{pq}{16|\tau_{n}|^{2}}(\frac{\nabla_{\bm{x}_{0}}p}{p}+\frac{\nabla_{\bm{x}_{0}}q}{q}-\frac{2}{|\tau_{n}|}\nabla_{\bm{x}_{0}}|\tau_{n}|)$
+$\nabla_{\bm{x}_{0}}\mu_{n}\nobreakspace=\frac{pq}{16|\tau_{n}|^{2}}(\frac{\nabla_{\bm{x}_{0}}p}{p}+\frac{\nabla_{\bm{x}_{0}}q}{q}-\frac{2}{|\tau_{n}|}\nabla_{\bm{x}_{0}}|\tau_{n}|)$
 where
@@
 For simplicity, we denote
-$c=\frac{1}{|\tau_{n}|},c_{0}=\frac{1}{pl_{0}}+\frac{1}{l_{0}^{2}},c_{1}=~\frac{1}{pl_{1}}+\frac{1}{l_{1}^{2}},c_{2}=\frac{1}{pl_{2}}+\frac{1}{l_{2}^{2}},$
+$c=\frac{1}{|\tau_{n}|},c_{0}=\frac{1}{pl_{0}}+\frac{1}{l_{0}^{2}},c_{1}=\nobreakspace\frac{1}{pl_{1}}+\frac{1}{l_{1}^{2}},c_{2}=\frac{1}{pl_{2}}+\frac{1}{l_{2}^{2}},$
 $c_{01}=c_{0}+c_{1},c_{02}=c_{0}+c_{2},c_{12}=c_{1}+c_{2}$
@@
 Then we can write it in matrix form
-$\nabla_{\bm{V}_{n}}\mu_{n}~=~\begin{bmatrix}\bm{A}_{n}~&~\bm{B}_{n}\\ -\bm{B}_{n}~&~\bm{A}_{n}\end{bmatrix}~\bm{V}_{n}$
+$\nabla_{\bm{V}_{n}}\mu_{n}\nobreakspace=\nobreakspace\begin{bmatrix}\bm{A}_{n}\nobreakspace&\nobreakspace\bm{B}_{n}\\ -\bm{B}_{n}\nobreakspace&\nobreakspace\bm{A}_{n}\end{bmatrix}\nobreakspace\bm{V}_{n}$
 with
-$\bm{A}_{n}=\mu_{n}~\begin{pmatrix}c_{12}~&~-c_{2}~&~-c_{1}~\\ -c_{2}~&~c_{20}~&~-c_{0}~\\ -c_{1}~&~-c_{0}~&~c_{01}\end{pmatrix},\bm{B}_{n}~=~\mu_{n}\begin{pmatrix}0~&~-c~&~c~\\ c~&~0~&~-c~\\ -c~&~c~&~0\end{pmatrix}$
+$\bm{A}_{n}=\mu_{n}\nobreakspace\begin{pmatrix}c_{12}\nobreakspace&\nobreakspace-c_{2}\nobreakspace&\nobreakspace-c_{1}\nobreakspace\\ -c_{2}\nobreakspace&\nobreakspace c_{20}\nobreakspace&\nobreakspace-c_{0}\nobreakspace\\ -c_{1}\nobreakspace&\nobreakspace-c_{0}\nobreakspace&\nobreakspace c_{01}\end{pmatrix},\bm{B}_{n}\nobreakspace=\nobreakspace\mu_{n}\begin{pmatrix}0\nobreakspace&\nobreakspace-c\nobreakspace&\nobreakspace c\nobreakspace\\ c\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-c\nobreakspace\\ -c\nobreakspace&\nobreakspace c\nobreakspace&\nobreakspace 0\end{pmatrix}$
 Matrix $\bm{A}_{n}$ is Laplacian and $\bm{B}_{n}$ is antisymmetric.
@@
 of the triangular mesh, we have
-$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}~&~\bm{B}~\\ -\bm{B}~&~\bm{A}~\\ \end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\end{bmatrix}=\bm{G}_{F}\bm{V}$
+$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}\nobreakspace&\nobreakspace\bm{B}\nobreakspace\\ -\bm{B}\nobreakspace&\nobreakspace\bm{A}\nobreakspace\\ \end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\end{bmatrix}=\bm{G}_{F}\bm{V}$
 where $\bm{A}$ is $N_{v}\times N_{v}$ Laplacian matrix, and $\bm{B}$ is $N_{v}\times N_{v}$ antisymmetric matrix,
@@
 The radius of inside sphere of $\tau_{n}$ is
-$r_{n}~=~\frac{3|\tau_{n}|}{s}$
+$r_{n}\nobreakspace=\nobreakspace\frac{3|\tau_{n}|}{s}$
 with $s=s_{0}+s_{1}+s_{2}+s_{3}$. The radius ratio metric of $\tau_{n}$ is
@@
 where
-$\nabla_{\bm{x}_{0}}|\bm{d}_{0}|=$ $\frac{1}{|\bm{d}_{0}|}\{2[\bm{d}_{0}\cdot(\bm{v}_{20}\times\bm{v}_{30})\bm{v}_{10}+\bm{d}_{0}\cdot(\bm{v}_{30}\times\bm{v}_{10})\bm{v}_{20}+\bm{d}_{0}\cdot(\bm{v}_{10}\times\bm{v}_{20})\bm{v}_{30}]~$
+$\nabla_{\bm{x}_{0}}|\bm{d}_{0}|=$ $\frac{1}{|\bm{d}_{0}|}\{2[\bm{d}_{0}\cdot(\bm{v}_{20}\times\bm{v}_{30})\bm{v}_{10}+\bm{d}_{0}\cdot(\bm{v}_{30}\times\bm{v}_{10})\bm{v}_{20}+\bm{d}_{0}\cdot(\bm{v}_{10}\times\bm{v}_{20})\bm{v}_{30}]\nobreakspace$
 $+\bm{d}_{0}\times[(\bm{v}_{30}^{2}-\bm{v}_{20}^{2})\bm{v}_{10}+(\bm{v}_{10}^{2}-\bm{v}_{30}^{2})\bm{v}_{20}+(\bm{v}_{20}^{2}-\bm{v}_{10}^{2})\bm{v}_{30}]\}$
@@
 $k_{23}=\bm{v}_{30}^{2}-\bm{v}_{20}^{2},k_{31}=\bm{v}_{10}^{2}-\bm{v}_{30}^{2},k_{12}=\bm{v}_{20}^{2}-\bm{v}_{10}^{2}$
-$[\bm{d}_{0}]_{\times}=\bm{D}_{0}=\begin{bmatrix}0~&~-d_{2}~&~d_{1}~\\ d_{2}~&~0~&~-d_{0}~\\ -d_{1}~&~d_{0}~&~0\end{bmatrix}$
+$[\bm{d}_{0}]_{\times}=\bm{D}_{0}=\begin{bmatrix}0\nobreakspace&\nobreakspace-d_{2}\nobreakspace&\nobreakspace d_{1}\nobreakspace\\ d_{2}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-d_{0}\nobreakspace\\ -d_{1}\nobreakspace&\nobreakspace d_{0}\nobreakspace&\nobreakspace 0\end{bmatrix}$
 so we have
@@
 Then we can write $\nabla_{\bm{V}_{n}}|\bm{d}_{0}|$ in matrix form
-$\nabla_{\bm{V}_{n}}|\bm{d}_{0}|~=~\frac{1}{12|\tau_{n}|R_{n}}\begin{bmatrix}\bm{M}~&~-d_{2}\bm{K}~&~d_{1}\bm{K}~\\ d_{2}\bm{K}~&~\bm{M}~&~-d_{0}\bm{K}~\\ -d_{1}\bm{K}~&~d_{0}\bm{K}~&~\bm{M}\end{bmatrix}~\bm{V}_{n}$
+$\nabla_{\bm{V}_{n}}|\bm{d}_{0}|\nobreakspace=\nobreakspace\frac{1}{12|\tau_{n}|R_{n}}\begin{bmatrix}\bm{M}\nobreakspace&\nobreakspace-d_{2}\bm{K}\nobreakspace&\nobreakspace d_{1}\bm{K}\nobreakspace\\ d_{2}\bm{K}\nobreakspace&\nobreakspace\bm{M}\nobreakspace&\nobreakspace-d_{0}\bm{K}\nobreakspace\\ -d_{1}\bm{K}\nobreakspace&\nobreakspace d_{0}\bm{K}\nobreakspace&\nobreakspace\bm{M}\end{bmatrix}\nobreakspace\bm{V}_{n}$
 with
-$\bm{M}=\begin{bmatrix}2c~&~-2c_{23}~&~-2c_{31}~&~-2c_{12}~\\ -2c_{23}~&~2c_{23}~&~0~&~0~\\ -2c_{31}~&~0~&~2c_{31}~&~0~\\ -2c_{12}~&~0~&~0~&~2c_{12}\end{bmatrix},\bm{K}=\begin{bmatrix}0~&~-k_{23}~&~-k_{31}~&~-k_{12}~\\ k_{23}~&~0~&~-\bm{v}_{30}^{2}~&~\bm{v}_{20}^{2}~\\ k_{31}~&~\bm{v}_{30}^{2}~&~0~&~-\bm{v}_{10}^{2}~\\ k_{12}~&~-\bm{v}_{20}^{2}~&~\bm{v}_{10}^{2}~&~0\end{bmatrix}$
+$\bm{M}=\begin{bmatrix}2c\nobreakspace&\nobreakspace-2c_{23}\nobreakspace&\nobreakspace-2c_{31}\nobreakspace&\nobreakspace-2c_{12}\nobreakspace\\ -2c_{23}\nobreakspace&\nobreakspace 2c_{23}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{12}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{12}\end{bmatrix},\bm{K}=\begin{bmatrix}0\nobreakspace&\nobreakspace-k_{23}\nobreakspace&\nobreakspace-k_{31}\nobreakspace&\nobreakspace-k_{12}\nobreakspace\\ k_{23}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-\bm{v}_{30}^{2}\nobreakspace&\nobreakspace\bm{v}_{20}^{2}\nobreakspace\\ k_{31}\nobreakspace&\nobreakspace\bm{v}_{30}^{2}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-\bm{v}_{10}^{2}\nobreakspace\\ k_{12}\nobreakspace&\nobreakspace-\bm{v}_{20}^{2}\nobreakspace&\nobreakspace\bm{v}_{10}^{2}\nobreakspace&\nobreakspace 0\end{bmatrix}$
 $c=c_{23}+c_{31}+c_{12}$, matrix $\bm{M}$ is symmetric, $\bm{K}$ is antisymmetric.
 For $\nabla_{\bm{x}}s$, we denote
-$p_{0}$ $=~\frac{\bm{v}_{31}^{2}}{4s_{2}}+\frac{\bm{v}_{21}^{2}}{4s_{3}}+\frac{\bm{v}_{32}^{2}}{4s_{1}},\quad~p_{1}~=~\frac{\bm{v}_{32}^{2}}{4s_{0}}+\frac{\bm{v}_{02}^{2}}{4s_{3}}+\frac{\bm{v}_{30}^{2}}{4s_{2}},\quad$
-$~p_{2}$ $=~\frac{\bm{v}_{30}^{2}}{4s_{1}}+\frac{\bm{v}_{10}^{2}}{4s_{3}}+\frac{\bm{v}_{31}^{2}}{4s_{0}},\quad p_{3}~=~\frac{\bm{v}_{10}^{2}}{4s_{2}}+\frac{\bm{v}_{20}^{2}}{4s_{1}}+\frac{\bm{v}_{21}^{2}}{4s_{0}},\quad$
+$p_{0}$ $=\nobreakspace\frac{\bm{v}_{31}^{2}}{4s_{2}}+\frac{\bm{v}_{21}^{2}}{4s_{3}}+\frac{\bm{v}_{32}^{2}}{4s_{1}},\quad\nobreakspace p_{1}\nobreakspace=\nobreakspace\frac{\bm{v}_{32}^{2}}{4s_{0}}+\frac{\bm{v}_{02}^{2}}{4s_{3}}+\frac{\bm{v}_{30}^{2}}{4s_{2}},\quad$
+$\nobreakspace p_{2}$ $=\nobreakspace\frac{\bm{v}_{30}^{2}}{4s_{1}}+\frac{\bm{v}_{10}^{2}}{4s_{3}}+\frac{\bm{v}_{31}^{2}}{4s_{0}},\quad p_{3}\nobreakspace=\nobreakspace\frac{\bm{v}_{10}^{2}}{4s_{2}}+\frac{\bm{v}_{20}^{2}}{4s_{1}}+\frac{\bm{v}_{21}^{2}}{4s_{0}},\quad$
 $q_{01}$ $=-(\frac{\bm{v}_{31}\cdot\bm{v}_{30}}{4s_{2}}+\frac{\bm{v}_{21}\cdot\bm{v}_{20}}{4s_{3}}),q_{02}=-(\frac{\bm{v}_{32}\cdot\bm{v}_{30}}{4s_{1}}+\frac{\bm{v}_{12}\cdot\bm{v}_{10}}{4s_{3}}),q_{03}=-(\frac{\bm{v}_{23}\cdot\bm{v}_{20}}{4s_{1}}+\frac{\bm{v}_{13}\cdot\bm{v}_{10}}{4s_{2}}),$
@@
 Then we can write $\nabla_{\bm{V}_{n}}s$ in matrix form
-$\nabla_{\bm{V}_{n}}\mu_{n}~=~\begin{bmatrix}\bm{S}~&~~&~~\\ ~&~\bm{S}~&~~\\ ~&~~&~\bm{S}\end{bmatrix}~\bm{V}_{n}$
+$\nabla_{\bm{V}_{n}}\mu_{n}\nobreakspace=\nobreakspace\begin{bmatrix}\bm{S}\nobreakspace&\nobreakspace\nobreakspace&\nobreakspace\nobreakspace\\ \nobreakspace&\nobreakspace\bm{S}\nobreakspace&\nobreakspace\nobreakspace\\ \nobreakspace&\nobreakspace\nobreakspace&\nobreakspace\bm{S}\end{bmatrix}\nobreakspace\bm{V}_{n}$
 with
-$\bm{S}~=~\begin{bmatrix}p_{0}&q_{01}&q_{02}&q_{03}\\ q_{01}&p_{1}&~q_{12}&~q_{13}\\ q_{02}&~q_{12}&p_{2}&q_{23}\\ q_{03}&q_{13}&q_{23}&p_{3}\end{bmatrix}$
+$\bm{S}\nobreakspace=\nobreakspace\begin{bmatrix}p_{0}&q_{01}&q_{02}&q_{03}\\ q_{01}&p_{1}&\nobreakspace q_{12}&\nobreakspace q_{13}\\ q_{02}&\nobreakspace q_{12}&p_{2}&q_{23}\\ q_{03}&q_{13}&q_{23}&p_{3}\end{bmatrix}$
 Matrix $\bm{S}$ is symmetric.
@@
 We denote
-$\bm{C}_{0}=\begin{bmatrix}0~&~x_{2}~&~x_{3}~&~x_{1}~\\ x_{3}~&~0~&~x_{0}~&~x_{2}~\\ x_{1}~&~x_{3}~&~0~&~x_{0}~\\ x_{2}~&~x_{0}~&~x_{1}~&0\end{bmatrix},\bm{C}_{1}=\begin{bmatrix}0~&~y_{2}~&~y_{3}~&~y_{1}~\\ y_{3}~&~0~&~y_{0}~&~y_{2}~\\ y_{1}~&~y_{3}~&~0~&~y_{0}~\\ y_{2}~&~y_{0}~&~y_{1}~&0\end{bmatrix},\bm{C}_{2}=\begin{bmatrix}0~&~z_{2}~&~z_{3}~&~z_{1}~\\ z_{3}~&~0~&~z_{0}~&~z_{2}~\\ z_{1}~&~z_{3}~&~0~&~z_{i}~\\ z_{2}~&~z_{0}~&~z_{1}~&0\end{bmatrix}$
+$\bm{C}_{0}=\begin{bmatrix}0\nobreakspace&\nobreakspace x_{2}\nobreakspace&\nobreakspace x_{3}\nobreakspace&\nobreakspace x_{1}\nobreakspace\\ x_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace x_{0}\nobreakspace&\nobreakspace x_{2}\nobreakspace\\ x_{1}\nobreakspace&\nobreakspace x_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace x_{0}\nobreakspace\\ x_{2}\nobreakspace&\nobreakspace x_{0}\nobreakspace&\nobreakspace x_{1}\nobreakspace&0\end{bmatrix},\bm{C}_{1}=\begin{bmatrix}0\nobreakspace&\nobreakspace y_{2}\nobreakspace&\nobreakspace y_{3}\nobreakspace&\nobreakspace y_{1}\nobreakspace\\ y_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace y_{0}\nobreakspace&\nobreakspace y_{2}\nobreakspace\\ y_{1}\nobreakspace&\nobreakspace y_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace y_{0}\nobreakspace\\ y_{2}\nobreakspace&\nobreakspace y_{0}\nobreakspace&\nobreakspace y_{1}\nobreakspace&0\end{bmatrix},\bm{C}_{2}=\begin{bmatrix}0\nobreakspace&\nobreakspace z_{2}\nobreakspace&\nobreakspace z_{3}\nobreakspace&\nobreakspace z_{1}\nobreakspace\\ z_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace z_{0}\nobreakspace&\nobreakspace z_{2}\nobreakspace\\ z_{1}\nobreakspace&\nobreakspace z_{3}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace z_{i}\nobreakspace\\ z_{2}\nobreakspace&\nobreakspace z_{0}\nobreakspace&\nobreakspace z_{1}\nobreakspace&0\end{bmatrix}$
 Then we have
-$\nabla_{\bm{V}_{n}}|\tau_{n}|=\frac{1}{6}\begin{bmatrix}0~&~-\bm{C}_{2}~&~\bm{C}_{1}~\\ \bm{C}_{2}~&~0~&~-\bm{C}_{0}~\\ -\bm{C}_{1}~&~\bm{C}_{0}~&~0\end{bmatrix}\bm{V}_{n}=\bm{C}\bm{V}_{n}$
+$\nabla_{\bm{V}_{n}}|\tau_{n}|=\frac{1}{6}\begin{bmatrix}0\nobreakspace&\nobreakspace-\bm{C}_{2}\nobreakspace&\nobreakspace\bm{C}_{1}\nobreakspace\\ \bm{C}_{2}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace-\bm{C}_{0}\nobreakspace\\ -\bm{C}_{1}\nobreakspace&\nobreakspace\bm{C}_{0}\nobreakspace&\nobreakspace 0\end{bmatrix}\bm{V}_{n}=\bm{C}\bm{V}_{n}$
 Where $|\tau_{n}|$ is a cubic form with respect to $\bm{V}_{n}$, we have the following lemma for cubic forms.
@@
 If the function $f$ is a cubic form with respect to $\bm{X}=(x_{0},x_{1},\cdots,x_{n})$, then there exists a third-order matrix $\bm{U}$ such that
-$f~=~U_{ijk}x_{i}x_{j}x_{k}$
+$f\nobreakspace=\nobreakspace U_{ijk}x_{i}x_{j}x_{k}$
 Let $W_{ik}^{j}=\frac{1}{2}(U_{ijk}+U_{kji})$, for any fixed $j$, $\bm{W}^{j}$ is a symmetric matrix, and we have
-$f~=W_{ik}^{j}~x_{i}~x_{j}x_{k}$
-So we have $\nabla_{x_{k}}f=U_{ijk}x_{i}x_{j}=W_{ik}^{j}x_{i}x_{j}$, let $C_{ki}~=~U_{ijk}x_{j},~E_{ki}~=~W_{ik}^{j}x_{j}$, so $\bm{E}~=~\frac{1}{2}(\bm{C}+\bm{C}^{T})$, then we can express $\nabla f$ as
+$f\nobreakspace=W_{ik}^{j}\nobreakspace x_{i}\nobreakspace x_{j}x_{k}$
+So we have $\nabla_{x_{k}}f=U_{ijk}x_{i}x_{j}=W_{ik}^{j}x_{i}x_{j}$, let $C_{ki}\nobreakspace=\nobreakspace U_{ijk}x_{j},\nobreakspace E_{ki}\nobreakspace=\nobreakspace W_{ik}^{j}x_{j}$, so $\bm{E}\nobreakspace=\nobreakspace\frac{1}{2}(\bm{C}+\bm{C}^{T})$, then we can express $\nabla f$ as
 $\nabla f=\bm{C}\bm{X}=\bm{E}\bm{X}$
 From the above theorem, we can get $\nabla_{\bm{V}_{n}}|\tau_{n}|=\frac{1}{2}(\bm{C}+\bm{C}^{T})\bm{V}_{n}$.
-$\frac{1}{2}(\bm{C}+\bm{C}^{T})=\frac{1}{12}\begin{bmatrix}0~&~\bm{C}_{2}^{T}-\bm{C}_{2}~&~\bm{C}_{1}^{T}-\bm{C}_{1}~\\ \bm{C}_{2}-\bm{C}_{2}^{T}~&~0~&~\bm{C}_{0}^{T}-\bm{C}_{0}~\\ \bm{C}_{1}^{T}-\bm{C}_{1}~&~\bm{C}_{0}^{T}-\bm{C}_{0}~&~0\end{bmatrix}$
+$\frac{1}{2}(\bm{C}+\bm{C}^{T})=\frac{1}{12}\begin{bmatrix}0\nobreakspace&\nobreakspace\bm{C}_{2}^{T}-\bm{C}_{2}\nobreakspace&\nobreakspace\bm{C}_{1}^{T}-\bm{C}_{1}\nobreakspace\\ \bm{C}_{2}-\bm{C}_{2}^{T}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace\bm{C}_{0}^{T}-\bm{C}_{0}\nobreakspace\\ \bm{C}_{1}^{T}-\bm{C}_{1}\nobreakspace&\nobreakspace\bm{C}_{0}^{T}-\bm{C}_{0}\nobreakspace&\nobreakspace 0\end{bmatrix}$
 It is evident that each block matrix in $\frac{1}{2}(\bm{C}+\bm{C}^{T})$ is an antisymmetric matrix.
@@
 and the gradients $\nabla_{\bm{V}_{n}}|\bm{d}_{0}|$, $\nabla_{\bm{V}_{n}}s$, and $\nabla_{\bm{V}_{n}}|\tau_{n}|$, we can obtain
-$\nabla_{\bm{V}_{n}}\mu_{n}=\mu_{n}\begin{bmatrix}\bm{A}_{n}~&~\bm{B}_{0n}~&~\bm{B}_{1n}~\\ -\bm{B}_{2n}~&~\bm{A}_{n}~&~\bm{B}_{0n}~\\ -\bm{B}_{1n}~&~-\bm{B}_{0n}~&~\bm{A}_{n}\end{bmatrix}\bm{V}_{n}$
+$\nabla_{\bm{V}_{n}}\mu_{n}=\mu_{n}\begin{bmatrix}\bm{A}_{n}\nobreakspace&\nobreakspace\bm{B}_{0n}\nobreakspace&\nobreakspace\bm{B}_{1n}\nobreakspace\\ -\bm{B}_{2n}\nobreakspace&\nobreakspace\bm{A}_{n}\nobreakspace&\nobreakspace\bm{B}_{0n}\nobreakspace\\ -\bm{B}_{1n}\nobreakspace&\nobreakspace-\bm{B}_{0n}\nobreakspace&\nobreakspace\bm{A}_{n}\end{bmatrix}\bm{V}_{n}$
 with
-$\bm{A}_{n}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}+\frac{1}{s}\bm{S},~\bm{B}_{0n}=\frac{-d_{0}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{0}^{T}-\bm{C}_{0}),$
+$\bm{A}_{n}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}+\frac{1}{s}\bm{S},\nobreakspace\bm{B}_{0n}=\frac{-d_{0}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{0}^{T}-\bm{C}_{0}),$
 $\bm{B}_{1n}=\frac{d_{1}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{1}^{T}-\bm{C}_{1}),\bm{B}_{2n}=\frac{-d_{2}}{12|\tau_{n}|R_{n}}\bm{K}-\frac{1}{6|\tau_{n}|}(\bm{C}_{2}^{T}-\bm{C}_{2})$
@@
 of the tetrahedral mesh,we have
-$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}~&~\bm{B}_{2}~&~\bm{B}_{1}~\\ -\bm{B}_{2}~&~\bm{A}~&~\bm{B}_{0}~\\ -\bm{B}_{1}~&~-\bm{B}_{0}~&~\bm{A}\end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\\ \bm{Z}\end{bmatrix}=\bm{G}_{F}\bm{V}~$
+$\nabla_{\bm{V}}F=\begin{bmatrix}\bm{A}\nobreakspace&\nobreakspace\bm{B}_{2}\nobreakspace&\nobreakspace\bm{B}_{1}\nobreakspace\\ -\bm{B}_{2}\nobreakspace&\nobreakspace\bm{A}\nobreakspace&\nobreakspace\bm{B}_{0}\nobreakspace\\ -\bm{B}_{1}\nobreakspace&\nobreakspace-\bm{B}_{0}\nobreakspace&\nobreakspace\bm{A}\end{bmatrix}\begin{bmatrix}\bm{X}\\ \bm{Y}\\ \bm{Z}\end{bmatrix}=\bm{G}_{F}\bm{V}\nobreakspace$
 where $\bm{A}$ is $N_{v}\times N_{v}$ symmetric matrix, and $\bm{B}_{0}$, $\bm{B}_{1}$, $\bm{B}_{2}$ are $N_{v}\times N_{v}$ antisymmetric matrices.
@@
 $\bm{S}$ is laplacian,
-$\bm{M}=\begin{bmatrix}2c~&~-2c_{23}~&~-2c_{31}~&~-2c_{12}~\\ -2c_{23}~&~2c_{23}~&~0~&~0~\\ -2c_{31}~&~0~&~2c_{31}~&~0~\\ -2c_{12}~&~0~&~0~&~2c_{12}\end{bmatrix}$
+$\bm{M}=\begin{bmatrix}2c\nobreakspace&\nobreakspace-2c_{23}\nobreakspace&\nobreakspace-2c_{31}\nobreakspace&\nobreakspace-2c_{12}\nobreakspace\\ -2c_{23}\nobreakspace&\nobreakspace 2c_{23}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{31}\nobreakspace&\nobreakspace 0\nobreakspace\\ -2c_{12}\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2c_{12}\end{bmatrix}$
 $c=c_{12}+c_{23}+c_{31}$. However, since the terms $c_{12},c_{23},c_{31}$ may take negative values, we define
-$\bm{M}^{abs}=\begin{bmatrix}2|c_{12}|+|c_{23}|+|c_{31}|~&~-2|c_{23}|~&~-2|c_{31}|~&~-2|c_{12}|~\\ -2|c_{23}|~&~2|c_{23}|~&~0~&~0~\\ -2|c_{31}|~&~0~&~2|c_{31}|~&~0~\\ -2|c_{12}|~&~0~&~0~&~2|c_{12}|\end{bmatrix}$
+$\bm{M}^{abs}=\begin{bmatrix}2|c_{12}|+|c_{23}|+|c_{31}|\nobreakspace&\nobreakspace-2|c_{23}|\nobreakspace&\nobreakspace-2|c_{31}|\nobreakspace&\nobreakspace-2|c_{12}|\nobreakspace\\ -2|c_{23}|\nobreakspace&\nobreakspace 2|c_{23}|\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace\\ -2|c_{31}|\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2|c_{31}|\nobreakspace&\nobreakspace 0\nobreakspace\\ -2|c_{12}|\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 0\nobreakspace&\nobreakspace 2|c_{12}|\end{bmatrix}$
 $\bm{A}_{n}^{abs}=\frac{1}{12|\tau_{n}|R_{n}}\bm{M}^{abs}+\frac{1}{s}\bm{S}$
@@
 Given the nodal coordinate vector $\bm{V}$ and the gradient matrix
-$\bm{G}_{F}=\begin{bmatrix}\bm{A}~&~\bm{B}_{2}~&~\bm{B}_{1}~\\ -\bm{B}_{2}~&~\bm{A}~&~\bm{B}_{0}~\\ -\bm{B}_{1}~&~-\bm{B}_{0}~&~\bm{A}\end{bmatrix}$
+$\bm{G}_{F}=\begin{bmatrix}\bm{A}\nobreakspace&\nobreakspace\bm{B}_{2}\nobreakspace&\nobreakspace\bm{B}_{1}\nobreakspace\\ -\bm{B}_{2}\nobreakspace&\nobreakspace\bm{A}\nobreakspace&\nobreakspace\bm{B}_{0}\nobreakspace\\ -\bm{B}_{1}\nobreakspace&\nobreakspace-\bm{B}_{0}\nobreakspace&\nobreakspace\bm{A}\end{bmatrix}$
 the optimization can be solved using the fixed-point iteration method.
 The gradient condition $\bm{G}_{F}\bm{V}=0$ can be formulated as a system of equations
-$\begin{cases}\bm{A}\bm{X}+\bm{B}_{2}\bm{Y}+\bm{B}_{1}\bm{Z}=\bm{0}~\\ -\bm{B}_{2}\bm{X}+\bm{A}\bm{Y}+\bm{B}_{0}\bm{Z}=\bm{0}~\\ -\bm{B}_{1}\bm{X}-\bm{B}_{0}\bm{Y}+\bm{A}\bm{Z}=\bm{0}\end{cases}$
+$\begin{cases}\bm{A}\bm{X}+\bm{B}_{2}\bm{Y}+\bm{B}_{1}\bm{Z}=\bm{0}\nobreakspace\\ -\bm{B}_{2}\bm{X}+\bm{A}\bm{Y}+\bm{B}_{0}\bm{Z}=\bm{0}\nobreakspace\\ -\bm{B}_{1}\bm{X}-\bm{B}_{0}\bm{Y}+\bm{A}\bm{Z}=\bm{0}\end{cases}$
 By employing the fixed-point iteration method for optimization with fixed boundaries and decomposing $\bm{V}=\bm{V}_{free}+\bm{V}_{fix}$, the system of equations can be expressed as
-$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}_{2}\bm{Y}^{old}-\bm{B}_{1}\bm{Z}^{old}-\bm{A}\bm{X}^{old}_{fix}~\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}_{2}\bm{X}^{old}-\bm{B}_{0}\bm{Z}^{old}-\bm{A}\bm{Y}^{old}_{fix}~\\ \bm{A}\bm{Z}_{free}^{new}=\bm{B}_{1}\bm{X}^{old}+\bm{B}_{0}\bm{Y}^{old}-\bm{A}\bm{Z}^{old}_{fix}\end{cases}$
+$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}_{2}\bm{Y}^{old}-\bm{B}_{1}\bm{Z}^{old}-\bm{A}\bm{X}^{old}_{fix}\nobreakspace\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}_{2}\bm{X}^{old}-\bm{B}_{0}\bm{Z}^{old}-\bm{A}\bm{Y}^{old}_{fix}\nobreakspace\\ \bm{A}\bm{Z}_{free}^{new}=\bm{B}_{1}\bm{X}^{old}+\bm{B}_{0}\bm{Y}^{old}-\bm{A}\bm{Z}^{old}_{fix}\end{cases}$
 In practical implementations, substituting $\bm{A}$ with $\bm{A}_{abs}$ achieves enhanced computational performance.
 For the two-dimensional case, the system of equations is formulated as
-$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}\bm{Y}^{old}-\bm{A}\bm{X}^{old}_{fix}~\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}\bm{X}^{old}-\bm{A}\bm{Y}^{old}_{fix}\end{cases}$
+$\begin{cases}\bm{A}\bm{X}_{free}^{new}=-\bm{B}\bm{Y}^{old}-\bm{A}\bm{X}^{old}_{fix}\nobreakspace\\ \bm{A}\bm{Y}_{free}^{new}=\bm{B}\bm{X}^{old}-\bm{A}\bm{Y}^{old}_{fix}\end{cases}$
 Let $d=\bm{V}^{new}-\bm{V}^{old}$ denote the search direction, and update the mesh nodal coordinate vector as
@@
 First, we initialize
-$\delta=\begin{cases}0,\quad\text{if}~{}k\leq~m~\\ k-m,\quad\text{if}~{}k>m\end{cases};L=\begin{cases}k,\quad\text{if}~{}k\leq~m~\\ m,\quad\text{if}~{}k>m\end{cases};\bm{q}_{L}=\nabla F(\bm{x}_{k})$
+$\delta=\begin{cases}0,\quad\text{if}~k\leq\nobreakspace m\nobreakspace\\ k-m,\quad\text{if}~k>m\end{cases};L=\begin{cases}k,\quad\text{if}~k\leq\nobreakspace m\nobreakspace\\ m,\quad\text{if}~k>m\end{cases};\bm{q}_{L}=\nabla F(\bm{x}_{k})$
 Each iteration of the PLBFGS algorithm is as follows:
@@
 13:end for
-14:  $d_{k}=\bm{r}_{L}$
-15:  $\bm{x}_{k+1}=\bm{x}_{k}+\lambda_{k}d_{k}$
+[…]
+[…]
 The matrix $\bm{P}_{k}$ is a chosen preconditioner, $\lambda_{k}$ is given by the line search satisfying the strong Wolfe conditions [32] with an initial guess $\lambda_{k}=1$, and
@@
 4: Update search direction $\bm{p}_{k+1}=-\bm{g}_{k+1}+\beta_{k+1}\bm{p}_{k}$
-$\lambda_{k}$ is given by line search satisfying the strong Wolfe conditions with an initial guess $\lambda_{k}=1$.
-The scalar coefficient $\beta_{k}$ has multiple calculation formulas, see [16].
-Notable $\beta_{k}$ formulas include: Polak-Ribière ($\beta_{k}^{PR}$[35], Hestenes–Stiefel ($\beta_{k}^{HS}$[18], Fletcher–Reeves ($\beta_{k}^{FR})$ [13], and Dai-Yuan($\beta_{k}^{DY}$[9].
-We adopt the Polak-Ribière $\beta_{k}$ formula, defined as:
+[…]
 $\beta_{k+1}^{PR}=\frac{\bm{g}_{k+1}^{T}(\bm{g}_{k+1}-\bm{g}_{k})}{\bm{g}_{k}^{T}\bm{g}_{k}}$
@@
 The radius ratio $\mu=\frac{R}{dr}$ of inscribed to circumscribed spheres is adopted as the quality metric.
-We run all experiments on a laptop running Ubuntu 22.04. The CPU is an AMD Ryzen7 4800H(2.9GHz,8MB cache) and 16GB DDR4 RAM(3200MHz).
+We run all experiments on a laptop running Ubuntu 22.04. The CPU is an AMD Ryzen7 7735H(3.2GHz,16MB cache) and 16GB DDR5 RAM(5600MHz).
 4.1 2D TriangleMesh
@@
 The results demonstrate that preconditioning significantly reduces both iteration counts and search steps(often by nearly half or more) and takes the less time, while achieving comparable or superior optimization outcomes to non-preconditioned cases.
-Figure 5: For the L-shaped domain, the corner points and points on the domain boundary edges are fixed, while points on the boundary faces are allowed to move along the tangential directions of the faces.
+Figure 5: This domain is an L-shaped domain with an empty ball domain inside.
+For the L-shaped domain, the corner points and points on the domain boundary edges are fixed, while points on the boundary faces are allowed to move along the tangential directions of the faces.
 ODT optimization converges after 36 iterations.
@@
 Sphere Init / 0.01317 391 19784 /
-ODT 15 0.05079 106 19072 4.18
-ODT 75 0.06376 58 18852 20.31
-LBFGS 17(85) 0.10295 53 19784 7.95
-PLBFGS 13(54) 0.18480 44 19784 5.84
-NLCG 30(107) 0.16996 45 19784 9.89
-PNLCG 16(69) 0.17120 43 19784 6.97
-ODT+LBFGS 15+15(59) 0.28190 1 19072 4.18+3.64
-ODT+PLBFGS 15+7(29) 0.30972 0 19072 4.18+2.16
-ODT+PNLCG 15+7(38) 0.29942 1 19072 4.18+2.65
-ODT+LBFGS 75+13(33) 0.31686 0 18852 20.31+2.04
-ODT+PLBFGS 75+7(27) 0.32546 0 18852 20.31+2.01
-ODT+PNLCG 75+8(34) 0.32572 0 18852 20.31+2.45
+ODT 15 0.05079 106 19072 2.70
+ODT 75 0.06376 58 18852 13.05
+LBFGS 29(110) 0.177782 42 19784 7.60
+PLBFGS 14(60) 0.192391 43 19784 4.82
+NLCG 28(125) 0.166748 41 19784 8.84
+PNLCG 13(57) 0.181469 42 19784 4.36
+ODT+LBFGS 15+15(59) 0.29190 1 19072 2.70+3.64
+ODT+PLBFGS 15+7(29) 0.30972 0 19072 2.70+2.06
+ODT+NLCG 15+15(62) 0.303464 0 19072 2.70+3.72
+ODT+PNLCG 15+7(34) 0.30249 1 19072 2.70+2.33
+ODT+LBFGS 75+13(33) 0.31686 0 18852 13.05+1.67
+ODT+PLBFGS 75+7(25) 0.32531 0 18852 13.05+1.52
+ODT+NLCG 75+14(56) 0.29968 0 18852 13.05+2.80
+ODT+PNLCG 75+8(37) 0.312975 0 18852 13.05+2.12
 Lshape Init / 0.00971 610 32827 /
-ODT 15 0.01938 225 31846 13.48
-ODT 36 0.06550 96 31371 31.49
-LBFGS 14(109) 0.00556 6 32827 13.49
-PLBFGS 24(145) 0.20427 10 32827 27.34
-NLCG 10(82) 0.01216 12 32827 14.06
-PNLCG 16(117) 0.12828 5 32827 22.22
-ODT+LBFGS 15+22(91) 0.24053 4 31846 13.48+12.07
-ODT+PLBFGS 15+12(61) 0.22503 4 31846 13.48+9.26
-ODT+PNLCG 15+14(76) 0.19577 4 31846 13.48+11.50
-ODT+LBFGS 36+12(51) 0.29676 1 31371 22.22+6.72
-ODT+PLBFGS 36+8(32) 0.34373 0 31371 22.22+5.08
-ODT+PNLCG 36+9(51) 0.31098 0 31371 22.22+7.57
+ODT 15 0.01938 225 31846 9.39
+ODT 36 0.06550 96 31371 21.65
+LBFGS 34(158) 0.22334 5 32827 20.78
+PLBFGS 25(112) 0.24498 6 32827 17.08
+NLCG 33(136) 0.209923 8 32827 17.69
+PNLCG 18(81) 0.23201 4 32827 12.34
+ODT+LBFGS 15+24(108) 0.21932 5 31846 9.39+10.67
+ODT+PLBFGS 15+11(48) 0.23547 4 31846 9.39+5.86
+ODT+NLCG 15+24(104) 0.20628 5 31846 9.39+10.38
+ODT+PNLCG 15+16(83) 0.21572 4 31846 9.39+10.16
+ODT+LBFGS 36+12(51) 0.29676 1 31371 21.65+5.03
+ODT+PLBFGS 36+8(32) 0.34373 0 31371 21.65+3.94
+ODT+NLCG 36+15(71) 0.34499 0 31371 21.65+6.95
+ODT+PNLCG 36+8(37) 0.35575 0 31371 22.22+4.58
 12 spheres
@@
 init / 0.00408 1470 70105 /
-ODT 15 0.01119 843 68962 30.35
-ODT 75 0.08138 188 66581 144.77
-LBFGS 48(224) 0.17128 215 70105 83.62
-PLBFGS 25(139) 0.19374 219 70105 56.42
-NLCG 37(188) 0.09946 235 70105 36.26
-PNLCG 16(88) 0.14231 217 70105 36.26
-ODT+LBFGS 15+21(107) 0.22678 78 68962 30.35+30.85
-ODT+PLBFGS 15+12(54) 0.21974 68 68962 30.35+17.92
-ODT+PNLCG 15+11(61) 0.23623 66 68692 30.35+19.89
-ODT+LBFGS 75+13(56) 0.24514 3 66581 144.77+15.74
-ODT+PLBFGS 75+6(10) 0.28008 2 66581 144.77+4.39
-ODT+PNLCG 75+7(35) 0.25638 3 66581 144.77+11.7
+ODT 15 0.01119 843 68962 20.14
+ODT 75 0.08138 188 66581 97.46
+LBFGS 42(192) 0.17838 220 70105 51.05
+PLBFGS 28(153) 0.17877 216 70105 45.28
+NLCG 33(148) 0.142602 219 70105 39.67
+PNLCG 15(68) 0.16193 216 70105 20.53
+ODT+LBFGS 15+25(112) 0.21412 70 68962 20.14+24.33
+ODT+PLBFGS 15+22(103) 0.23569 68 68962 20.14+24.98
+ODT+NLCG 15+24(113) 0.20821 69 68692 20.14+23.65
+ODT+PNLCG 15+11(53) 0.20433 66 68692 20.14+12.89
+ODT+LBFGS 75+13(56) 0.24514 3 66581 97.46+15.74
+ODT+PLBFGS 75+6(10) 0.28008 2 66581 97.46+3.37
+ODT+NLCG 75+14(59) 0.21657 5 66581 97.46+11.91
+ODT+PNLCG 75+7(28) 0.28628 3 66581 97.46+6.74
 Table  1 presents quality statistics and efficiency comparisons for all mesh models, with iteration counts in parentheses indicating Wolfe line search steps.
@@
 We also design an efficient preconditioner to substantially enhance optimization efficiency.
-Our algorithm can operate independently or synergistically with existing mesh optimization methods.
+Our algorithm can operate independently or synergistically with existing mesh optimization methods, such as ODT algorithm.
 Benefiting from the preconditioner, it significantly improves mesh quality with minimal computational overhead when applied as a post-processing step to other optimization techniques.
-Our algorithm is a vertex smoothing algorithm that does not include topological optimization.
-It performs well when mesh topology is already good.
-When our algorithm is combined with ODT algorithm, it can achieve good optimization results because ODT algorithm includes topology changes and effectively optimizes the mesh topology.
-When used alone, our algorithm can also optimize sliver elements, although the results are less effective than when combined with ODT.
-Furthermore, for surface boundaries, we currently use a fixed boundary point method, which also limits the optimization effect.
-Therefore, in future works, we plan to develop more effective boundary processing solution and add an edge/face exchange algorithm to further improve the algorithm, enhance the reliability of the algorithm, and extend this mesh optimization algorithm to anisotropic mesh.
+In future works, we plan to develop more effective boundary processing solution and to further improve the algorithm, and extend this mesh optimization algorithm to anisotropic mesh.
 We also hope to optimize larger-scale mesh and further improve computational efficiency by using GPU parallel technology.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
 Acknowledgement